below. (20 pts.) ar A 10.8x Industria de ovs s Fast Retailing Co Ltd Ltd LLP SA Median Maxi 31.12.2017 225.0
Answer #1

Answer:

Mean EV/EBITDA multiple = 14.3

MAVI's EBITDA = 225.0 million

Hence:

MAVI's Enterprise Value (EV) = 225.0 * 14.3 = $3,217.50 million

Net Debt = 221.5

Value of Equity = 3,217.50 - 221.5 = 2,996 million

Number of shares = 49,657,000

Stock price = 2,996 million / 49,657,000 = 60.33

TRY (million) EBITDA 31.12.2017 225.00 Million 51.80 Million 221.50 Million Net debt EV/EBITDA Number of shares Stock price 14.3 49,657,000 60.33